Solve each system by addition, if possible
5x - 2y = 19
3x + 4y = 1

Answer by stanbon(75887)

Solve each system by addition, if possible
5x - 2y = 19
3x + 4y = 1
------
Modify for elimination::
10x - 4y = 38
3x + 4y = 1
------
Add and solve for "x":
13x = 39
x = 3
-----
Solve for "y":
5x - 2y = 19
5*3 - 2y = 19
15 - 2y = 19
-2y = 4
y = -2
-------------
Cheers,
Stan H.
================